A 2.78 g sample of hydrated iron (II) sulfate was heated to remove all water of hydration. The mass of the anhydrous compound was 1.52 g. Determine the molecular formula of the original hydrated compound 


1
Expert's answer
2021-01-11T03:50:32-0500

The concept is very simple. Hydrates are compounds that contain a certain amount of water. When you heat a hydrate, all the water that's a part of its structure is evaporated, leaving behind the anhydrous salt.

In your case, the hydrated form of iron (II) sulfate weighs 2.78 g. After all the water of hydration is removed, you're left with the anhydrous salt iron (II) sulfate, which weighs 1.52 g.

The difference in weight between the hydrate and the anhydrous salt is the water of hydration.

mwater = mhydrate - manhydrous salt

mwater = 2.78 - 1.52 = 1.26 g


To determine the number of water molecules associated with each unit of iron (II) sulfate, you need to use the two compounds' molar masses

1.26g ("\frac{1molewater}{18.015g}") = 0.0700mole water


1.52g ("\frac{1moleFeSO_4}{151.91g}") = 0.0100mole FeSO4


Divide these numbers by the smallest one to get the mole ratio that exists between water and iron (II) sulfate in the hydrate


H2O = 0.0700/0.0100 = 7 mole

FeSO4 = 0.0100/0.0100 = 1 mole


This means that each unit of hydrated iron (II) sulfate contains 7 moles of water


FeSO4 • 7 H2O iron (II) sulfate heptahydrate